# Learning in Markets with Heterogeneous Agents: Dynamics and Survival of Bayesian vs. No-Regret Learners

### v1

@@ 2 Model and Preliminaries
We use the market structure as in [17] in our analysis of wealth and price dynamics. At each time in discrete steps indexed by $t$, a state is realized. There is a finite number of states $s\in\{1,\dots,S\}$ distributed i.i.d. with probability $q=(q_{1},\dots,q_{S})$ where $q_{s}\geq\Delta>0$ for all $s$. At each time, there is a complete set of assets. Assets are also indexed by $s$ with one unit of asset $s$ purchased at time $t$ paying one unit of wealth at time $t+1$ if state $s$ occurs at time $t+1$ and 00 otherwise. These assets, known as Arrow securities, are more general than they may seem; any asset structure that permits arbitrary reallocation of wealth across states can be written as a simple transformation of Arrow securities. This asset market model, originally introduced in [4], is the standard general equilibrium model of an economy with a complete set of assets, see [3] Chapter 19. Asset prices at time $t$ are represented by $p_{t}=(p_{1t},\dots,p_{St})$.
There are $N$ agents indexed by $n=1,\dots,N$. The wealth of agent $n$ at time $t$ is $w^{n}_{t}$. Agents begin with initial wealths $w^{n}_{0}>0$ for agent $n$. These initial wealths are normalized to sum to one. [^3] This normalization is done only for convenience. We analyze relative wealths. The actual size of the economy (the aggregate wealth) does not play a role in relative wealth dynamics and may follow an arbitrary stochastic process, as long as it is bounded away from zero and bounded from above. Thus, if state $s$ occurs at time $t+1$, agent $n$ will have time $t+1$ wealth, $w^{n}_{t+1}=\frac{\alpha^{n}_{st}w^{n}_{t}}{p_{st}}$.
@@
$w^{n}_{t+1}=\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{p_{st}}\bigg{)}^{\mathbf{1}_{\{s_{t}=s\}}}w^{n}_{t},$ (2)
where $\mathbf{1}_{\{s_{t}=s\}}$ is the indicator function taking on the value $1$ if the state realized at time $t$ is state $s$, i.e., $s_{t}=s$, and equals zero otherwise. Equivalently, the wealth at time $t>0$ is
$w^{n}_{t}=\prod_{\tau=1}^{t}\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{s\tau}}{p_{s\tau}}\bigg{)}^{\mathbf{1}_{\{s_{\tau}=s\}}}w^{n}_{0}.$ (3)
@@
$r^{nm}_{t+1}$ $=\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}}\bigg{)}^{\mathbf{1}_{\{s_{t}=s\}}}r^{nm}_{t},$ (4)
$r^{nm}_{T}$ $=\prod_{t=1}^{T}\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}}\bigg{)}^{\mathbf{1}_{\{s_{t}=s\}}}r^{nm}_{0}.$ (5)
@@
$\log(r^{nm}_{T})=\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}})+\log(r^{nm}_{0}).$ (6)
@@
An agent $n$ survives in the market if there exists a constant market share $c$ such that
@@
Let $\{\alpha^{n}_{t}\}_{t=0}^{\infty}$ be a sequence of investment profiles for agent $n$. If there exists a atate $s\in S$ for which $\alpha^{n}_{st}=0$ infinitely often, then agent $n$ vanishes from the market.
This straightforward observation immediately rules out certain learning approaches, such as those that treat asset selection as pure actions. It demonstrates that any reasonable learning strategy must be restricted to investment profiles with full support on the assets. Specifically, all learning strategies we consider will assign at least a small, constant probability $\Delta$ to each state.
@@ 3 Comparing Learners
As we noted above, an investor who knows the probabilities $q_{s}$ maximizes the expected growth rate of their wealth at each time $t$ using the investment rule $\alpha^{n}_{t}=q$. Next, consider an agent $n$ with a sequence of investment profiles $\alpha_{t}^{n}$. We will first compare the wealth of this agent to that of an agent who uses the investment rule $\alpha=q$, and then, we will compare different types of learning agents.
As we saw in the previous section in Equation (6), the log of the wealth ratio of an agent at time $T$ compared to an investor using the optimal $\alpha=q$ rule can be expressed as
$\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log\alpha^{n}_{st}-\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log q_{s}+\log(r^{nm}_{0}).$ (10)
The main goal of this paper is to compare different theories on how to best learn to invest. One theory is Bayesian learning, assuming that the investor $n$ has a Bayesian prior over the possible asset return probabilities. At each step, a Bayesian learner uses the expected return probability implied by their beliefs, updating their beliefs according to the observed states by using Bayes rule. One can also use multi-armed bandit learning methods given such a prior of possible asset return distributions, and have the agent invest using one of the distributions in its prior at each step, aiming to learn which one is the best. A different theory would consider the maximization problem of Equation (10) as a concave maximization problem and use no-regret learning algorithms to optimize the value.
From the latter two perspectives, if the difference between the two sums in Equation (10) is negative, this is a form of regret [^5] This is regret only with respect to $q$; while $q$ is optimal in expectation, it may not be the best fixed strategy in hindsight for a given sequence of state realizations. We extend on this point later; see Theorem 5.2. against a fixed strategy $\alpha=q$. First, consider an investor with a fixed number of possible models for $q$, $\{\theta^{1},\ldots,\theta^{K}\}$. This formulation of the problem would allow us to consider an investor using a classical multi-arm bandit approach, such as UCB (see, e.g., [75]): the investor can invest using one of the $\theta^{k}$ as investment rules, and use no-regret methods to update their belief on the expected value of each of the options. Assuming the correct distribution $q$ is among the options considered, we can express the gap in value between any one choice $\theta^{k}$ and the true distribution using Equation (9) as
@@
In the next section, we consider a Bayesian investor who not only has a finite list of possible $q$ vectors, but also has a prior over these models, and uses Bayes rule to update their prior given the outcomes seen in previous rounds.
@@
$U^{t}(\alpha)=\sum_{s}\mathbf{1}_{\{s_{t}=s\}}\log\alpha_{st}.$ (13)
@@
Note that the regret of $\alpha^{n}_{1:T}$ is at least as large as its regret against the investor with the optimal fixed rule $\alpha=q$. The maximum fixed rule $\alpha$ used in the definition is the best fixed investment rule with hindsight, having seen the outcome of all choices and random state realizations, which is not an implementable investment rule even for an investor who knows the probabilities. So even the investor with knowledge of the probabilities and using $\alpha^{n}_{t}=q$ for all $t$ will have some regret based on the outcome of random events (see Section 5 and Theorem 5.2).
@@
$\log(r^{nm}_{T})=\sum_{t=1}^{T}U^{t}(\alpha^{n}_{t})-\sum_{t=1}^{T}U^{t}(\alpha^{m}_{t})+\log(r^{nm}_{0})=R^{T}(\alpha_{1:T}^{m})-R^{T}(\alpha_{1:T}^{n})+\log(r^{nm}_{0}).$ (16)
We note that despite this very strong benchmark for regret, and without any prior, the convex optimization literature offers very strong guarantees on expected regret. For any concave function $U^{t}(\alpha)$, a simple gradient-decent based algorithm guarantees regret at most $\sqrt{T}$ over $T$ periods, while for the $\log$ based utility function (that is Exp-Convex) considered here, second order optimization methods guarantee a $O(\log T)$ regret bound [15] (see also, [26] Theorem 3.3 or [56] Theorem 4.4.)
The central question we ask in this paper is: What learning algorithms are better in ensuring that the player’s wealth remains high, or at the very least, does not vanish? How do Bayesian learners and convex optimizers compare in terms of relative wealth in this game? Using equation (16) we get the following result, connecting regret and survival in the market.
@@
Proof.
This result follows from the above analysis, leading to Equation (16). For the first part, we have $R^{n}(T)-R^{m}(T)=-\log(r^{nm})-\log(r_{0}^{nm})=\log(r^{mn})+const$. If the left hand side diverges to +$\infty$+ ∞, then $r^{mn}=w^{m}/w^{n}$ diverges as well, which is only possible if $w^{n}\rightarrow 0$. For the second part, clearly this is a necessary condition due to the first part of the theorem: if an agent vanishes, it does not survive. To see that this condition is sufficient, if there exists a constant $c>0$ such that $R^{n}(T)-R^{m}(T)<c$ at all times for all $m\neq n$, then the wealth ratio $r^{mn}$ with any other player is bounded, and thus $w^{n}$ is bounded from below. ∎
@@ 4 Bayesian Learners
Next, we consider the standard Bayesian learning results. We restate and summarize in this section known results for completeness and to provide a framework for our comparison of Bayesian and no-regret learning.
Consider a Bayesian watching the process on states in the market. Suppose that the Bayesian does not know the probability $q$, and instead considers a finite number of possibilities for $q$: models $\theta^{1},\dots,\theta^{K}$. We assume that $\theta^{k}_{s}\geq\Delta$ for all $k$ and $s$. See the remark at the end of this section for connection to results for Bayesian learners with a continuum prior. The Bayesian has a prior $\lambda_{0}=(\lambda^{1}_{0},\dots,\lambda^{K}_{0})$ with $\lambda^{k}_{0}>0$ for all $k$. Let $n_{t}^{s}$ be the number of times that state $s$ occurs by time $t$. The posterior probability of model $\theta^{k}$ is, by Bayes rule,
@@
If model $k$ is correct, $\theta^{k}=q$, then $I_{q}(\theta^{k})=0$, so for any other model $\theta^{l}$ with $I_{q}(\theta^{l})\neq 0$ we have $\lambda^{l}_{t}\to 0$ and $\lambda^{k}_{t}\to 1$ almost surely. Equation (20) shows that the log of posterior probabilities converges linearly in time, and thus that the ratio of posterior probabilities converges at an exponential rate.
The Bayesian learns the true model if it is one of the models considered. Alternatively, suppose that no model considered by the Bayesian is correct, but that there is a unique model with minimum entropy relative to the true model, i.e., there is a $k^{*}$ such that $0<I_{q}(\theta^{k^{*}})<I_{q}(\theta^{k})$ for all $k\neq k^{*}$. Then it follows from equation (20) that $\lambda_{t}^{k^{*}}\to 1$ almost surely.
These convergence results can be summarized in the following proposition.
@@
Let $\Theta$ denote the support of a Bayesian learner’s prior, and let $\theta^{*}=\operatorname*{argmin}_{\theta\in\Theta}I_{q}(\theta)$. The Bayesian’s beliefs converge to $\theta^{*}$ almost surely.
@@ 4.1 An Investment Rule for the Bayesian
$s.t.$ $\sum_{s}\alpha^{n}_{st}=1,$
@@ 4.2 Wealth Dynamics for the Bayesian
A Bayesian agent that maximizes the expected growth rate of its wealth uses the investment rule $\alpha^{n}_{t}=q^{n}_{t}$ at each time $t$ (i.e., betting its beliefs). We consider the evolution of the wealth of the Bayesian relative to any other trader who uses an investment rule that does not depend on future states, i.e., it’s time $t$ value depends only on state realizations through time $t$. The log ratio of their wealths is, from Section 2
$\log(r^{nm}_{T})$ $=\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{n}_{st}}{q_{s}})-\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{m}_{st}}{q_{s}})+\log(r^{nm}_{0}).$ (23)
@@
$\sum_{t=1}^{T}\operatorname{\mathbb{E}}[\log(\frac{\alpha^{n}_{st}}{q_{s}})]-\operatorname{\mathbb{E}}[\log(\frac{\alpha^{m}_{st}}{q_{s}})]$ $=\sum_{t=1}^{T}I_{q}(\alpha^{m}_{t})-\sum_{t=1}^{T}I_{q}(\alpha^{n}_{t})+\log(r^{nm}_{0}).$ (24)
@@
Let $r^{nm}_{T}$ be the wealth ratio of a Bayesian learner $n$ competing against an agent $m$ playing $\alpha^{m}_{t}=q$ for all $t$. Then $\liminf_{T\rightarrow\infty}\operatorname{\mathbb{E}}[r^{nm}_{T}]>0$ almost surely; i.e., the Bayesian survives.

### v2

@@ 2 Model and Preliminaries
We use the market structure as in [17] in our analysis of wealth and price dynamics. At each time in discrete steps indexed by $t$, a state is realized. There is a finite number of states $s\in\{1,\dots,S\}$ distributed i.i.d. with probability $q=(q_{1},\dots,q_{S})$ where $q_{s}\geq\Delta>0$ for all $s$. At each time, there is a complete set of assets. Assets are also indexed by $s$ with one unit of asset $s$ purchased at time $t$ paying one unit of wealth at time $t+1$ if state $s$ occurs at time $t+1$ and 00 otherwise. These assets, known as Arrow securities, are more general than they may seem; any asset structure that permits arbitrary reallocation of wealth across states can be written as a simple transformation of Arrow securities. This asset market model, originally introduced by [3], is the standard general equilibrium model of an economy with a complete set of assets, see [76] Chapter 19. Asset prices at time $t$ are represented by $p_{t}=(p_{1t},\dots,p_{St})$.
There are $N$ agents indexed by $n=1,\dots,N$. The wealth of agent $n$ at time $t$ is $w^{n}_{t}$. Agents begin with initial wealths $w^{n}_{0}>0$ for agent $n$. These initial wealths are normalized to sum to one. [^3] This normalization is done only for convenience. We analyze relative wealths. The actual size of the economy (the aggregate wealth) does not play a role in relative wealth dynamics and may follow an arbitrary stochastic process, as long as it is bounded away from zero and bounded from above. Let $\alpha^{n}_{st}$ be the fraction of agent $n$’s wealth invested in asset $s$ at time $t$. Thus, if state $s$ occurs at time $t+1$, agent $n$ will have time $t+1$ wealth, $w^{n}_{t+1}=\frac{\alpha^{n}_{st}w^{n}_{t}}{p_{st}}$.
@@
$w^{n}_{t+1}=\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{p_{st}}\bigg{)}^{\mathbf{1}_{s_{t}=s}}w^{n}_{t},$ (2)
where $\mathbf{1}_{s_{t}=s}$ is the indicator function taking on the value $1$ if the state realized at time $t$ is state $s$, i.e., $s_{t}=s$, and equals zero otherwise. Equivalently, the wealth at time $t>0$ is
$w^{n}_{t}=\prod_{\tau=1}^{t}\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{s\tau}}{p_{s\tau}}\bigg{)}^{\mathbf{1}_{s_{\tau}=s}}w^{n}_{0}.$ (3)
@@
$r^{nm}_{t+1}$ $=\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}}\bigg{)}^{\mathbf{1}_{s_{t}=s}}r^{nm}_{t},$ (4)
$r^{nm}_{T}$ $=\prod_{t=1}^{T}\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}}\bigg{)}^{\mathbf{1}_{s_{t}=s}}r^{nm}_{0}.$ (5)
@@
$\log(r^{nm}_{T})=\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{s_{t}=s}\log(\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}})+\log(r^{nm}_{0}).$ (6)
@@
An agent $n$ survives in the market if there exists a constant market share $c$ s.t.
@@
Let $\{\alpha^{n}_{t}\}_{t=0}^{\infty}$ be a sequence of investment profiles for agent $n$. If there exists a state $s\in S$ for which $\alpha^{n}_{st}=0$ infinitely often, then the agent $n$ vanishes from the market.
This simple observation immediately rules out certain learning approaches, such as those that treat asset selection as pure actions. It demonstrates that any reasonable learning strategy must be restricted to investment profiles with full support over the assets. Specifically, all learning strategies we consider will assign at least a small constant probability $\Delta$ to each state.
@@ 3 Comparing Learners
As we noted above, an investor who knows the probabilities $q_{s}$ maximizes the expected growth rate of their wealth at each time $t$ by using the investment rule $\alpha^{n}_{t}=q$. Next, consider an agent $n$ with a sequence of investment profiles $\alpha_{t}^{n}$. We will first compare the wealth of this agent with that of an agent who uses the investment rule $\alpha^{n}_{t}=q$, and then compare different types of learning agents.
As we saw in the previous section in Equation (6), the log of the wealth ratio of an agent at time $T$ compared to an investor using the optimal $\alpha^{n}_{t}=q$ rule can be expressed as
$\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{s_{t}=s}\log\alpha^{n}_{st}-\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{s_{t}=s}\log q_{s}+\log r^{nm}_{0}.$ (10)
The main goal of this paper is to compare different theories on how to best learn to invest. One theory is Bayesian learning, assuming that the investor $n$ has a Bayesian prior over the possible asset return probabilities. At each step, a Bayesian learner uses the expected return probability implied by their beliefs, updating their beliefs according to the observed states using Bayes’ rule. One can also use multi-armed bandit learning methods given such a prior of possible asset return distributions, and have the agent invest using one of the distributions in its prior at each step, aiming to learn which one is the best. A different theory would consider the maximization problem of Equation (10) as a concave maximization problem and use no-regret learning algorithms to optimize the value.
From the latter two perspectives, if the difference between the two sums in Equation (10) is negative, this is a form of regret [^5] This is regret only with respect to $q$; while $q$ is optimal in expectation, it may not be the best fixed strategy in hindsight for a given sequence of state realizations. We expand on this point later; see Theorem 5.2. against a fixed strategy $\alpha=q$. First, consider an investor with a fixed number of possible models for $q$, $\{\theta^{1},\ldots,\theta^{K}\}$. This formulation of the problem would allow us to consider an investor using a classical multi-arm bandit approach, such as UCB (see, e.g., [85]): the investor can invest using one of the $\theta^{k}$ as investment rules, and use no-regret methods to update their belief on the expected value of each of the options. Assuming the correct distribution $q$ is among the options considered, we can express the gap in value between any one choice $\theta^{k}$ and the true distribution using Equation (9) as
@@
In the next section, we consider a Bayesian investor who not only has a finite list of possible $q$ vectors, but also has a prior over these models, and uses Bayes’ rule to update their prior given the outcomes seen in previous rounds.
@@
$U^{t}(\alpha)=\sum_{s}\mathbf{1}_{s_{t}=s}\log\alpha_{st}.$ (13)
@@
Note that the regret of $\alpha^{n}_{1:T}$ is at least as large as its regret against the investor with the optimal fixed rule $\alpha=q$. The maximum fixed rule $\alpha$ used in the definition is the best fixed investment rule with hindsight, having seen the outcome of all choices and random state realizations, which is not an implementable investment rule even for an investor who knows the probabilities. So even the investor with knowledge of the probabilities and using $\alpha^{n}_{t}=q$ for all $t$ will have some regret based on the outcome of random events (see Section 5 and Theorem 5.2 for a detailed analysis).
@@
$\log(r^{nm}_{T})$ $=\sum_{t=1}^{T}U^{t}(\alpha^{n}_{t})-\sum_{t=1}^{T}U^{t}(\alpha^{m}_{t})+\log(r^{nm}_{0})$ (16)
$=R^{T}(\alpha_{1:T}^{m})-R^{T}(\alpha_{1:T}^{n})+\log(r^{nm}_{0}).$
So, given the initial wealth shares, the agents’ regret levels specify their resulting wealths.
We note that despite this very strong benchmark for regret, and without any prior, the convex optimization literature offers very strong guarantees on expected regret. For any concave function $U^{t}(\alpha)$, a simple gradient-descent based algorithm guarantees regret at most $\sqrt{T}$ over $T$ periods, while for the $\log$-based utility function (that is Exp-Convex) considered here, second-order optimization methods guarantee a $O(\log T)$ regret bound [15] (see also [27] Theorem 3.3 or [61] Theorem 4.4.)
The central question we ask is: What learning algorithms are better in ensuring that the player’s wealth remains high, or at the very least, does not vanish? How do Bayesian learners and no-regret convex optimizers compare in terms of relative wealth? Using Equation (16), we get the following result, connecting regret and survival in the market. The proof is given in Appendix A.
@@ 4 Bayesian Learners
Next, we consider the standard Bayesian learning analysis. We restate and summarize in this section known results for completeness and to provide a framework for our comparison of Bayesian and no-regret learning.
Consider a Bayesian watching the process on states in the market. Suppose that the Bayesian does not know the probability $q$, and instead considers a finite number of possibilities for $q$: models $\theta^{1},\dots,\theta^{K}$. We assume that $\theta^{k}_{s}\geq\Delta$ for all $k$ and $s$. See the remark at the end of this section for connection to results on Bayesian learners with a continuum prior. The Bayesian has a prior $\lambda_{0}=(\lambda^{1}_{0},\dots,\lambda^{K}_{0})$ with $\lambda^{k}_{0}>0$ for all $k$. Let $n_{t}^{s}$ be the number of times that state $s$ occurs by time $t$. The posterior probability of model $\theta^{k}$ is, by Bayes rule,
@@
If model $k$ is correct, $\theta^{k}=q$, then $I_{q}(\theta^{k})=0$, so for any other model $\theta^{l}$ with $I_{q}(\theta^{l})\neq 0$ we have $\lambda^{l}_{t}\to 0$ and $\lambda^{k}_{t}\to 1$ almost surely as $t\rightarrow\infty$. Equation (20) shows that the log of posterior probabilities converges linearly in time, and thus, that the ratio of posterior probabilities converges at an exponential rate.
The Bayesian learns the true model if it is one of the models considered. Alternatively, suppose that no model considered by the Bayesian is correct, but that there is a unique model with minimum entropy relative to the true model; i.e., there exists $k^{*}$ such that $0<I_{q}(\theta^{k^{*}})<I_{q}(\theta^{k})$ for all $k\neq k^{*}$. Then it follows from equation (20) that $\lambda_{t}^{k^{*}}\to 1$ almost surely as $t\rightarrow\infty$. These convergence results can be summarized in the following proposition.
@@
Let $\Theta$ denote the support of a Bayesian learner’s prior, and let $\theta^{*}=\operatorname*{argmin}_{\theta\in\Theta}I_{q}(\theta)$. The Bayesian’s posterior probability on $\theta^{*}$ converges to $1$ almost surely.
@@ 4.1 An Investment Rule for the Bayesian
$s.t.$ $\sum_{s}\alpha^{n}_{st}=1.$
Using the definition of wealth, we have
@@ 4.2 Wealth Dynamics for the Bayesian
A Bayesian agent that maximizes the expected growth rate of its wealth uses the investment rule $\alpha^{n}_{t}=q^{n}_{t}$ at each time $t$ (i.e., betting its beliefs). We consider the evolution of the wealth of the Bayesian relative to any other trader $m$ who uses an investment rule that does not depend on future states, i.e., its time-$t$ value depends only on state realizations through time $t$. The log ratio of their wealths is, from Section 2,
$\log(r^{nm}_{T})$ $=\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{s_{t}=s}\log(\frac{\alpha^{n}_{st}}{q_{s}})$ (23)
$-\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{s_{t}=s}\log(\frac{\alpha^{m}_{st}}{q_{s}})+\log(r^{nm}_{0}).$
@@
$\sum_{t=1}^{T}$ $\operatorname{\mathbb{E}}[\log(\frac{\alpha^{n}_{st}}{q_{s}})]-\operatorname{\mathbb{E}}[\log(\frac{\alpha^{m}_{st}}{q_{s}})]$ (24)
$=\sum_{t=1}^{T}I_{q}(\alpha^{m}_{t})-\sum_{t=1}^{T}I_{q}(\alpha^{n}_{t})+\log(r^{nm}_{0}).$
@@
Let $r^{nm}_{T}$ be the wealth ratio of a Bayesian learner indexed by $n$ competing against an agent indexed by $m$ playing $\alpha^{m}_{t}=q$ for all $t$. Then $\liminf_{T\rightarrow\infty}\operatorname{\mathbb{E}}[r^{nm}_{T}]>0$ almost surely; i.e., the Bayesian survives.